OpenStudy (anonymous):

Which operation should be performed first according to the order of operations? 44 ÷ 2 + 9 – [120 – (5 + 10 • 7)] A. 44 ÷ 2 B. 2 + 9 C. 5 + 20 D. 10 • 7

OpenStudy (anonymous):

Here it is

OpenStudy (anonymous):

do what is in the inner parentheses first

OpenStudy (anonymous):

sorry my internet is soo slow.

OpenStudy (anonymous):

um c.....

OpenStudy (anonymous):

\[44\div 2+9-(120-\color{red}{(5+10\times 7)})\]

OpenStudy (anonymous):

105

OpenStudy (astrophysics):

How does it go again, bedmas is it, so you're close, try again

OpenStudy (anonymous):

the inner most parentheses are in red look there first

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

i.e. look only ast \[\color{red}{5+10\times 7}\]what would you do first there?

OpenStudy (anonymous):

5+10

OpenStudy (anonymous):

mmm no

OpenStudy (anonymous):

10*7

OpenStudy (anonymous):

multiplications come before additions

OpenStudy (anonymous):

yes \(10\times 7\) first

OpenStudy (anonymous):

2 more left